A car horn has a frequency of 3500Hz. An observer hears the car horn approaching with a frequency of 3750Hz. Calculate the speed of the car in m/s. The speed of sound in the air is 343m/s.
1
Expert's answer
2020-05-20T09:27:26-0400


 The observer detects waves with a frequency f given by


"\nu=\nu_0\frac{s}{s-u}\to u=\frac{\nu-\nu_0}{\nu}s=\frac{3750-3500}{3750}\cdot343=22.87m/s"